OpenStudy (zachgod):

F is a cubic polynomial written as f(x)=ax^3+bx^2+cx+d find conditions on a,b,c such as F has (i) two critical points. (ii) one critical point. (iii) no critical points. Using the results from above to solve: (b) If F has only one critical point, then x=?

myininaya (myininaya):

So you know how to find critical numbers?

OpenStudy (zachgod):

All I have is y prime. y=3ax^2+2b+c than use guad formula on that.

myininaya (myininaya):

And when do you have 2 real solutions ?

myininaya (myininaya):

When what is bigger than 0?

myininaya (myininaya):

it starts with a d

myininaya (myininaya):

discr....

myininaya (myininaya):

and it is y=3ax^2+2bx+c do you know what d word I'm talking about?

OpenStudy (zachgod):

discrimination

myininaya (myininaya):

discriminant

myininaya (myininaya):

not discrimination that is what happens when people don't like another group of people

OpenStudy (zachgod):

Yea. My bad. sorry.

OpenStudy (zachgod):

Im totally lost.

myininaya (myininaya):

so recall if you have y=ax^2+bx+c two real solutions will happen if b^2-4ac>0 one real solution will happen if b^2-4ac=0 no real solutions will happen if b^2-4ac<0

myininaya (myininaya):

So find the discriminant for your equation.

OpenStudy (zachgod):

Yep. I have that. But i don't know how to find the discriminant for the equation.

myininaya (myininaya):

\[y=Ax^2+Bx+C\] \[D=B^2-4AC\] Your equation is \[y'=3ax^2+2bx+c\] What is A for our equation?

OpenStudy (zachgod):

have no idea

myininaya (myininaya):

A is the thing in front of x^2 what is the thing in front of your x^2?

OpenStudy (zachgod):

3

myininaya (myininaya):

what about 3a is not that in front of x^2?

OpenStudy (zachgod):

Yea. 3a

OpenStudy (zachgod):

I didn't think it was important to write the A

myininaya (myininaya):

look we are comparing these equations: \[y=Ax^2+Bx+C\] \[y=3ax^2+2bx+c\] A=3a <--these were both the things in front of x^2 now what is B=?

OpenStudy (zachgod):

2b

myininaya (myininaya):

ok and course C=c so what is our discriminant D?

OpenStudy (zachgod):

d just goes away. i thnk

myininaya (myininaya):

\[D=B^2-4AC\] what is D with the values we just found for A,B, and C?

OpenStudy (zachgod):

no clue

myininaya (myininaya):

well you told B=2b and A=3a and C=c so you can find D b replacing those values in D

OpenStudy (zachgod):

is d=3

OpenStudy (zachgod):

is that the answer?

myininaya (myininaya):

No.